# City-on-Web: Real-time Neural Rendering of Large-scale Scenes on the Web

### v1

@@ 4.3 Optimization
We use Charbonnier loss [4] for reconstruction and S3IM loss [44] to assist multiple blocks’ model in capturing high-frequency details. Additionally, we use the interlevel loss to provide supervision signal for proposal MLP and distortion loss to reduce floaters like Mip-nerf 360 [1].
Moreover, we random uniform sample points set $\mathcal{P}$ within the bounding box of the scene and apply $L_{1}$ regularization on the alpha to encourage the blocks’ model to predict sparse occupied space:
${\mathcal{L}_{\textrm{sparse}}}=\frac{1}{|\mathcal{P}|}\sum_{p_{i}\in\mathcal{P}}|\alpha_{i}|=\frac{1}{|\mathcal{P}|}\sum_{p_{i}\in\mathcal{P}}|1-\sigma_{i}v|,$ (5)
where $v$ is the step size used in real-time rendering. Additionally, we introduce a regularization term for the opacity of the block. This regularization encourages the opacity of the block to be as close to 0 or 1 as possible, implying either full transparency or full opaqueness:
@@
In summary, the overall loss function is:
$\begin{split}{\mathcal{L}_{\textrm{train}}}=\;&{\mathcal{L}_{\textrm{charbonnier}}}+\lambda_{1}{\mathcal{L}_{\textrm{S3IM}}}+\lambda_{2}{\mathcal{L}_{\textrm{interlevel}}}+\\ &\lambda_{3}{\mathcal{L}_{\textrm{distortion}}}+\lambda_{4}{\mathcal{L}_{\textrm{sparse}}}+\lambda_{5}{\mathcal{L}_{\textrm{opacity}}}.\end{split}$ (7)
Method BlockA BlockE Campus Rubble Building
Metric
@@
[…]
NeRFacto
25.28
0.469
0.691
24.61
0.441
0.687
23.47
0.255
0.689
19.02
0.538
0.512
17.70
0.442
0.502
Instant-NGP
24.16
0.595
0.643
22.93
0.599
0.619
24.93
0.380
0.703
20.37
0.629
0.478
17.92
0.625
0.424
Mega-NeRF
25.24
0.566
0.668
25.61
0.468
0.680
22.28
0.472
0.565
23.68
0.558
0.525 20.30
0.526
0.506
Grid-NeRF 25.37
0.478
0.705
24.43
0.483
0.693
Ours 26.07 0.357 0.739 25.07 0.357 0.756 24.73 0.192 0.736 21.32 0.482 0.539 19.71 0.439 0.520
Table 1: Quantitative comparison on five large scene datasets. We report PSNR, LPIPS, and SSIM on the test views. The best and second best results are highlighted. The ’-’ symbol indicates that Grid-NeRF [45] was not evaluated on these datasets due to difficulties in adjusting its training configs beyond the provided configurations for Matrix City [19], resulting in poor performance on other datasets.
4.4 LOD Generation
To guarantee superior rendering quality from elevated perspectives and simultaneously diminish the resource demand for distant scene elements, our method involves generating multiple LOD for the scene. We successfully attained the scene’s lowest LOD in the training stage, ensuring maximal visual fidelity. For the downsampling and integration baking of multiple block models into a unified model, we initially freeze the training of hash encoding and decoder MLP components within these models. Subsequently, we proceed to retrain a new tiny deferred MLP. Following the deferred MLP’s successful retraining, we simulate lower resolution virtual voxels and triplane grid corners within the scenes of these multiple blocks. Lastly, this retrained deferred MLP is refined in collaboration with the network responsible for generating grid corner attributes, thereby optimizing the entire rendering process.
4.5 Baking
After the training stage, we conduct block-based evaluation and store the MLP’s outputs onto discrete grids, which generate segmented rendering resources. This approach facilitates efficient resource management for real-time rendering, as each block’s resources are handled independently. Initially, we render all training rays to collect ray samples. Samples with alpha and weight values above a certain threshold are retained, and samples below the threshold are discarded. The preserved samples are used to mark the adjacent eight grid points as occupied in the binary grids. After generating binary grids to identify occupied voxels, we follow the MERF by baking high-resolution 2D planes and a low-resolution 3D voxel grid in each block. Only the non-empty 3D voxels are stored using a block-sparse format. We downsample the occupancy grid with max-pooling for efficient rendering and skip empty space. To further save storage, we compress textures into the PNG format.
5 Experiments
Figure 4: Comparison to Mega-NeRF in Rubble dataset. The Rubble dataset presents significant variations in lighting. Although our method recovers more detail than Mega-NeRF, our deferred shading model has limited ability to represent view-dependent colors and cannot accurately represent lighting and exposure changes in the data as view-dependent effects. This limitation results in slightly lower PSNR values.
5.1 Experiments setup
Dataset and Metric. Our experiments span across various scales and environments. We have incorporated a real-world urban scene dataset (Campus) and public datasets consisting of real-world rural rubble scenes (Rublle, Building[37] and synthetic city-scale data (BlockA and BlockE in MatrixCity[19]. Our datasets were recorded under uniform, cloudy lighting conditions to minimize variation. To obtain precise pose information, we employed an annular capturing approach, which has a higher overlap rate compared to grid-based capturing methods. Tab. 2 presents an overview of our dataset. To assess the quality and fidelity of our reconstructions, we employ various evaluation metrics, including PSNR, SSIM and LPIPS [49].
Area$(m^{2})$ Images Overlap rate Resolution Altitude
1200$\times$800 6515 $95\%$ 8192$\times$5460 180m
Table 2: Overview of Campus dataset.
Implementations and Baselines. Our method takes posed multi-view images captured using a fly-through camera as input. The training code is built on the nerfstudio framework [35] with tiny-cuda-nn [25] extension. And our real-time viewer is a JavaScript web application whose rendering is implemented through GLS. We set the $512^{3}$ resolution for the voxel and $2048^{2}$ resolution for the triplane within each block. We use a 4-layer MLP with 64 hidden dimensions as an encoder after multi-resolution hash encoding to output density, color, and specular feature. Moreover, a 3-layer MLP with 16 hidden dimensions tiny deferred MLP is developed to predict residual view-dependent color. We sample $16384$ rays per batch and use Adam optimizer with an initial learning rate of $1\times 10^{-2}$ decaying exponentially to $1\times 10^{-3}$. Our model is trained with $50k$ iterations on one NVIDIA A100 GPU. We split the scene into 24 non-overlapping blocks for Campus scene and split other scenes into four blocks. Moreover, we benchmark current real-time rendering methods using three critical parameters: Payload (PL), GPU Memory (VRAM), and Frames Per Scond (FPS). Payload refers to the essential data transmitted during the rendering process. We perform qualitative comparisons between our method and existing SOTA methods for large-scale reconstruction. The Campus dataset is partitioned into six sections based on the reconstruction content. NeRFacto, Instant-NGP, and Grid-NeRF were applied to one of these sections, while in other datasets , they are applied to the entire scene. NeRFacto and Instant-NGP are utilized with the highest hash encoding resolution of $8192^{3}$. Similarly, Mega-NeRF divides the Campus dataset into 24 blocks and other datasets into four blocks. Our experiments focus on a single campus section for comparative analysis with existing real-time rendering methods.
5.2 Results Analysis
We systematically evaluate the performance of both baseline models and our method through qualitative and quantitative comparisons in Tab. 1 and Fig. 3. Notably, our method demonstrates a remarkable enhancement in visual fidelity as reflected by the SSIM and LPIPS metrics, which indicate the extent of detail restoration. Despite a reduction in PSNR compared to the SOTA methods, this is attributable to the fact that LPIPS and SSIM are more sensitive to the recovery of fine details, whereas PSNR mainly measures pixel-wise color accuracy. Our approach achieves higher fidelity reconstructions, revealing finer details due to our partitioned reconstruction strategy. In our method, each ray is evaluated by the deferred MLP only once, as opposed to other methods that evaluate the MLP at every sample point. Consequently, while our method recovers more intricate geometrical detail, it frequently results in color discrepancies with the ground truth image due to unstable lighting conditions and variable exposure, as shown in Fig. 4.
[…]
MobileNeRF
19.99
0.516
0.712
544.8
515.3
223
BakedSDF
22.24
0.627
0.413
712.3
242.1
68
Ours(Block)
24.82 0.741 0.190 128.1 40.6
51
Ours(Total)
526.6
114.4
46
Table 3: Comparison with existing real-time methods. We segment the scene into four blocks, while other methods reconstruct the whole scene. ‘Ours(Block)’ denotes the resource usage for one of these blocks, providing a fair comparison with other methods that process the scene as a single block. ‘Ours (Total)’ represents the cumulative resource usage for the entire scene.
In our evaluation, detailed in Tab. 3, we compare our method with current real-time rendering methods, using one segment of the Campus dataset for testing. These tests, are performed on an NVIDIA RTX 3060 Laptop GPU at a $1920\times 1080$ resolution. The results demonstrate that our method excels in reconstruction quality. We represent each scene block using voxels and triplanes, and store the baked grid attributes as images. This strategy significantly reduces the payload. This reduction notably accelerates resource transmission for web-based rendering applications. However, it is observed that our frame rate during rendering is lower compared to other methods. This is attributed to their rendering pipeline based on mesh rasterization, in contrast to our method, which utilizes volume rendering.
5.3 LOD Result
[…]
LOD3
23.72
0.660
0.306
132.1
40.2
LOD2
24.23
0.682
0.297
841.6
201.7
LOD1
24.73
0.736
0.192
3970.2
1259.6
Table 4: The LOD results on the whole Campus dataset.
Figure 5: Visualization of our LOD result.
Tab. 4 presents the quantitative rendering results at various LOD, along with the corresponding payload and VRAM usage. With increasing LOD, the resources required for rendering significantly decrease. Notably, our method’s lowest LOD level still maintains high fidelity rendering results, as demonstrated in Fig. 5. Our LOD strategy significantly streamlines the management of resource loading on web platforms, which is particularly advantageous in rendering distant blocks, as it requires less VRAM. It is worth noting that the VRAM usage presented in Tab. 4 represents the cumulative memory consumption of all blocks. Our dynamic loading strategy adaptively selects resources to load based on the camera’s field of view and the distance to each block, effectively keeping the peak VRAM usage around 1100MB.
5.4 Ablation Study
In Tab. 5, we conduct an ablation study of our method on one section of Campus dataset. Our model is trained for four blocks with low resolution (voxel resolution of $512^{3}$ and triplane resolution of $2048^{2}$). We also train a single model for the entire scene with high resolution (voxel resolution of $1024^{3}$ and triplane resolution of $4096^{2}$). Due to our non-overlapping scene partitioning strategy, these two representations have the same resolution across the entire scene. However, our reconstruction quality is higher, achieving better FPS and lower GPU memory usage, as well as reduced payload. We also removed our consistent training, including virtual grids and alpha blending during training. As shown in the table, our consistent training significantly improves reconstruction quality.
[…]
high res. model
24.05
0.710
0.201
540.9
147
32
w/o consistent
24.21
0.702
0.281
514.4
110
49
ours
24.82
0.741
0.190
526.6
114
46
Table 5: Ablation Study. The result is test on the one section of Campus dataset.
Figure 6: Limitations of Web-on-City.
6 Conclusion and Discussion
In this work, we introduced City-on-Web, which to our knowledge is the first system that enables real-time neural rendering of large-scale scenes over web using laptop GPUs. Our integration of block partitioning with LOD has significantly reduced the payload on the web platform and improved resource management efficiency. We ensured high-fidelity rendering quality by maintaining consistency between training and rendering. Extensive experiments have also fully proved the effectiveness of City-on-Web.
Limitation & Future Work. As shown in Fig. 6, our method still has some limitations. Since we derive alpha blending across shaders based on the Lambertian surface assumption, visible seams may occur at the boundaries between blocks on non-Lambertian surfaces, such as water surfaces. Combining physically-based rendering with multiple shaders blending may alleviate this problem. The deferred MLP in City-on-Web has limited representation ability for view-dependent color, which might cause numerous near-camera floaters. Taking the similar strategy used in [42, 12] by utilizing priors to pre-trim the scene or preprocessing the data are possible solutions.
Supplementary Material
7 Proof of 3D consistency
For a given sampling point $i$, suppose it is located within the region $\mathcal{K}$ of block $k$, which contains a total of $N_{k}$ sampling points. The diffuse color, feature, and opacity output by the shader of block $k$ are denoted as $\boldsymbol{c}^{k}_{i}$, $\boldsymbol{F}^{k}_{i}$, and $\alpha^{k}_{i}$ respectively. Then, for this ray, the output diffuse color $\boldsymbol{c}^{k}_{d}$ and specular feature $\boldsymbol{F}^{k}$of block $k$ are caculated as follows. $\boldsymbol{h}^{k}$ represents either the diffuse color or specular feature.
$\boldsymbol{h}^{k}(\boldsymbol{r})$ $=\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{h}_{i}^{k}$ (8)

### v2

@@ 4.3 Optimization
In the training stage, we reconstruct the finest LOD model by optimizing it with various losses:
$\mathcal{L}_{\text{train}}=\mathcal{L}_{\text{cb}}+\mathcal{L}_{\text{global}}+\lambda_{1}\mathcal{L}_{\text{s3im}}+\lambda_{2}\mathcal{L}_{\text{prop}}+\lambda_{3}\mathcal{L}_{\text{dist}}+\lambda_{4}\mathcal{L}_{\text{s}}+\lambda_{5}\mathcal{L}_{\text{opacity}}.$ (6)
Here, we use Charbonnier loss [3] $\mathcal{L}_{cb}$ for reconstruction and S3IM loss [44] $\mathcal{L}_{s3im}$ to assist model in capturing high-frequency details. Additionally, we use the interlevel loss $\mathcal{L}_{prop}$ to provide a supervision signal for proposal MLP and distortion loss $\mathcal{L}_{dist}$ to reduce floaters like Mip-nerf 360 [1].
Sparsity Loss. We random uniform sample points set $\mathcal{P}$ within the bounding box of the scene and apply $L_{1}$ regularization on the opacity of sample points $\alpha_{i}$ to encourage model to predict sparse occupied space:
${\mathcal{L}_{\textrm{sparse}}}=\frac{1}{|\mathcal{P}|}\sum_{p_{i}\in\mathcal{P}}|\alpha_{i}|$ (7)
Opacity Loss. We introduce a regularization term for the opacity of the block. This regularization encourages the opacity of the block to be as close to 0 or 1 as possible, implying either full transparency or full opaqueness:
@@
Regularization of Deferred MLPs. In the deferred rendering context, various combinations of specular and diffuse colors can satisfy multi-view consistency constraints. This situation often leads to incorrect disentanglement of these color components. Our training process involves using a deferred MLP within each block, but this approach does not guarantee the smoothness of specular color across block boundaries and the multitude of possible combinations of specular and diffuse colors.
Inspired by Grid-NeRF [45], which utilizes the smoothness of MLP to regularize explicit grid representations. We also utilize a global deferred MLP to regularize the rendering outputs from smaller, block-specific deferred MLPs, ensuring the global smoothness of specular color. In particular, we combine the specular color generated by this global deferred MLP with the diffuse color to obtain the final rendering result. We then supervise the rendering result using ground truth images in the form of a Charbonnier loss, denoted as $\mathcal{L}_{global}$, to regularize the smaller deferred MLPs. Notably, this global deferred MLP is significantly larger and thus possesses sufficient representational capacity compared to the smaller deferred MLPs designated for each block. Therefore, the global MLP does not limit the model’s representational capacity.
4.4 LOD Generation
To ensure high-quality rendering from elevated viewpoints and reduce resource usage for distant scene blocks, our method generates multiple LODs for the scene. One conventional approach to generate LODs would be to retrain the entire scene using fewer blocks, that is, at a lower representation resolution, but this method extends the training time a lot. Additionally, considering the specialized photography techniques employed for capturing large scenes, usually from aerial or top-down perspectives, it is challenging to ensure appearance consistency across models trained separately for extrapolated views if we retrain the entire scene from scratch.
Therefore, we generate LODs based on the scene’s finest LOD acquired during the training stage. Specifically, we simulate the virtual grid to store rendering attributes like MERF in the training stage. As merging $M\times M$ blocks into $\frac{M}{2}\times\frac{M}{2}$ blocks to generate LODs, we initially downsample the resolution of the virtual grid in each block by a factor of 2. Subsequently, we freeze the training of the query function $f^{k}$ within these submodels and retrain a new shared deferred MLP $\Phi^{k,l}$ across merged blocks. Finally, we continue to jointly optimize these submodels and the deferred MLP to adapt to lower-resolution voxels and triplanes.
4.5 Baking
For LOD $l$, we merge $M\times M$ blocks into $\frac{M}{2^{l}}\times\frac{M}{2^{l}}$ blocks. Every $2^{l}\times 2^{l}$ blocks can be baked into a single texture asset $f^{k,l}$ in the baking stage thanks to the dowmsampling when generating LODs. Thus, in the rendering stage, a single shader is responsible for rendering these $2^{l}\times 2^{l}$ blocks.
Specifically, we render all training rays to collect ray samples initially. Samples with opacity and weight values above a certain threshold are retained, and samples below the threshold are discarded. The preserved samples are used to mark the adjacent eight grid points as occupied in the binary occupancy grids $\phi^{k,l}$. After generating binary grids to identify occupied voxels, we follow MERF by baking high-resolution 2D planes and a low-resolution 3D voxel grid in each block to get the attribute function $f^{k,l}$ used in the rendering stage. Only the non-empty 3D voxels are stored using a block-sparse format. We downsample the occupancy grid with max-pooling for efficient rendering. To further save storage, we compress textures into the PNG format.
5 Experiments
Table 1: Quantitative comparison on the Matrix City, Campus, and Rubble datasets. We report PSNR, LPIPS, and SSIM on the test views. The best and second best results are highlighted.
Matrix City Campus Rubble
@@ 5 Experiments
NeRFacto 24.95 0.456 0.688 23.47 0.255 0.689 19.02 0.538 0.512
Instant-NGP 23.55 0.597 0.629 21.91 0.478 0.549 20.37 0.629 0.478
Mega-NeRF 25.43 0.517 0.674 22.28 0.472 0.565 23.68 0.558 0.525
Grid-NeRF 24.90 0.480 0.698 - - - - - -
Ours 25.87 0.332 0.734 24.73 0.192 0.736 21.32 0.482 0.539
5.1 Experiments setup
Dataset and Metric. Our experiments span across various scales and environments. We have incorporated a real-world urban scene dataset (Campus) and public datasets consisting of real-world rural rubble scenes (Rublle[39] and synthetic city-scale data (MatrixCity[17]. Our datasets were recorded under uniform, cloudy lighting conditions to minimize variation. To obtain precise pose information, we employed an annular capturing approach, which has a higher overlap rate compared to grid-based capturing methods. The dataset covers an area of $1200\times 800$ square meters. It includes a total of 6515 images. We use $99\%$ data for training, and the rest is used as the test dataset. To assess the quality and fidelity of our reconstructions, we employ various evaluation metrics, including PSNR, SSIM and LPIPS [49].
Implementations and Baselines. Our method takes posed multi-view images captured using a fly-through camera as input. The training code is built on the nerfstudio framework [36] with tiny-cuda-nn [25] extension. Our real-time viewer is a JavaScript web application whose rendering is implemented through GLSL. We set the $512^{3}$ resolution for the voxel and $2048^{2}$ resolution for the triplane within each block. We use a 4-layer MLP with 64 hidden dimensions as an encoder after multi-resolution hash encoding to output density, color, and specular feature. Moreover, a 3-layer MLP with 16 hidden dimensions tiny deferred MLP is developed to predict residual view-dependent color. We sample $16384$ rays per batch and use Adam optimizer with an initial learning rate of $1\times 10^{-2}$ decaying exponentially to $1\times 10^{-3}$. The global deferred MLP is a 6-layer MLP with 128 hidden dimensions. Our model is trained with $50k$ iterations on one NVIDIA A100 GPU. We split the scene into 24 blocks for Campus scene, and split other scenes into four blocks to reconstruct these scenes. We perform qualitative comparisons between our method and existing SOTA methods for large-scale reconstruction. The Campus dataset is partitioned into six parts. NeRFacto, Instant-NGP, and Grid-NeRF were applied to reconstruct one of these parts. NeRFacto and Instant-NGP are utilized with the highest hash encoding resolution of $8192^{3}$. Mega-NeRF divides the Campus scene into 24 blocks and splits another dataset into four blocks to evaluate its performance.
Our experiments focus on a single part in Campus dataset for comparative analysis with existing real-time rendering methods. This part contains over 1600 images, covering an area of approximately $600\times 400$ square meters. We divide this part of the data into four parts for a fair comparison with the MERF method. Other methods did not divide the dataset when conducting experiments. Moreover, we benchmark current real-time rendering methods using three critical parameters: Peak GPU memory usage (VRAM), frames per second (FPS), and on-disk storage (DISK). We report these metrics based on tests conducted on an NVIDIA RTX 3060 GPU with $1920\times 1080$ resolution.
5.2 Results Analysis
Figure 5: Visualization of our LOD result.
We systematically evaluate the performance of both baseline models and our method through qualitative and quantitative comparisons in Tab. 1 and Fig. 4. Notably, our method demonstrates a remarkable enhancement in visual fidelity as reflected by the SSIM and LPIPS metrics, which indicate the extent of detail restoration. Despite a reduction in PSNR compared to the SOTA methods, this is attributable to the fact that LPIPS and SSIM are more sensitive to the recovery of fine details, whereas PSNR mainly measures pixel-wise color accuracy. Our approach achieves higher fidelity reconstructions, revealing finer details due to our partitioned reconstruction strategy.
Table 2: Comparison with existing real-time methods. We divide the scene into four blocks and 16 blocks. We split the data into four parts, with each part being reconstructed by an individual MERF for a fair comparison. We present the results of Gauss splatting [16] after training 30,000 and 100,000 iterations for demonstration. VRAM and DISK are denoted in megabytes (MB).
@@ 5.2 Results Analysis
MobileNeRF 19.99 0.516 0.712 712.3 242.1 68
BakedSDF 22.24 0.627 0.413 544.8 515.3 223
MERF(4 blocks) 24.02 0.713 0.254 592.4 121.2 58
GS(3w iters) 23.78 0.745 0.263 1469.3 1469.1 77
GS(10w iters) 24.94 0.783 0.227 1467.1 1467.1 58
Ours(4 Blocks) 24.82 0.741 0.190 526.6 114.4 46
Ours(16 Blocks) 25.13 0.779 0.167 2040.7 464.5 34
In our evaluation, detailed in Tab. 2, we compare our method with current real-time rendering methods, using one part of the Campus dataset for testing. These tests, are performed on an NVIDIA RTX 3060 Laptop GPU at a $1920\times 1080$ resolution. The results demonstrate that our method excels in reconstruction quality. We represent each scene block using voxels and triplanes, and store the baked grid attributes as images. This strategy significantly reduces the memory. This reduction notably accelerates resource transmission for web-based rendering applications. However, it is observed that our frame rate during rendering is lower compared to other methods. This is attributed to their rendering pipeline based on mesh rasterization, which is in contrast to our method, which utilizes volume rendering.
5.3 LOD Result
Tab. 3 presents the quantitative rendering results at various LODs, along with the corresponding DISK and VRAM usage. With increasing LOD, the resources required for rendering significantly decrease. Notably, our method’s lowest LOD still maintains high-fidelity rendering results, as demonstrated in Fig. 5. Our LOD strategy significantly streamlines the management of resource loading on web platforms, which is particularly advantageous in rendering distant blocks, as it requires less VRAM. It is worth noting that the VRAM usage presented in Tab. 3 represents the cumulative memory consumption of all blocks. Our dynamic loading strategy adaptively selects resources to load based on the camera’s field of view and the distance to each block, effectively keeping the peak VRAM usage around 1100MB.
Table 3: The LOD results on the whole Campus dataset. VRAM and DISK are denoted in megabytes (MB).
@@ 5.3 LOD Result
LOD3 23.72 0.660 0.306 132.1 40.2
LOD2 24.23 0.682 0.297 841.6 201.7
LOD1 24.73 0.736 0.192 3970.2 1259.6
5.4 Ablation Study
We conduct ablation studies to demonstrate the impact of the contributions introduced to our method.
Ablation on Our Method. In Tab. 4, we conduct an ablation study of our method on one part of Campus dataset “ours(4 blocks)” means we use four blocks with $512^{3}$ voxel resolution and $2048^{3}$ triplane resolution for scene reconstruction. “ours(16 blocks)” means we use 16 blocks with $512^{3}$ voxel resolution and $2048^{3}$ triplane resolution for scene reconstruction. In “model with $1024^{3}$ Res.”, we train a one block MERF model with $1024^{3}$ voxel resolution and $4096^{2}$ triplane resolution. In “model with $2048^{3}$ Res.”, we train a one block MERF model with $2048^{3}$ voxel resolution and $8192^{2}$ triplane resolution. Our method has higher rendering quality and requires less storage space. In “no alpha blending”, we instead our alpha blending with simply using $\alpha_{i}/(\sum_{j}\alpha_{j})$ as blending weights. This non-occlusion-aware blending strategy significantly reduces the rendering quality. In “no consistent training”, we use MERF’s volume rendering pipeline in the training stage. In “no global deferred MLP”, we remove global deferred MLP. Without the regularization of deferred MLPs, the quality of the reconstruction has decreased.
Table 4: Ablation study on our method. The result is tested on one section of the Campus dataset. VRAM and DISK are denoted in megabytes (MB).
@@ 5.4 Ablation Study
model with $1024^{3}$ Res. 24.05 0.710 0.201 540.9 147
model with $2048^{3}$ Res. 24.42 0.751 0.184 3073.2 457
no alpha blending 24.03 0.684 0.345 565.4 153
no consistent training 24.21 0.702 0.281 514.4 110
no global deferred mlp 24.61 0.712 0.198 536.6 126
ours(4 blocks) 24.82 0.741 0.190 526.6 114
ours(16 Blocks) 25.13 0.779 0.167 2040.7 464
Ablation on LOD Generation. Tab. 5 shows ablation study on LOD generation. We use our LOD generation method as basline. In “downsample”, we simply downsample the model without re-optimization. In “retrain from scratch”, we do not use the finest LOD model to generate LOD. Instead, we trained the same resolution model from scratch.
Table 5: Ablation study on LOD generation.
@@
downsample 22.43 0.614 0.362 0hours
retrain from scratch 24.17 0.722 0.224 12hours
ours 24.20 0.724 0.204 4hours
6 Conclusion
In this work, we introduced City-on-Web, which to our knowledge is the first system that enables real-time neural rendering of large-scale scenes on the web using laptop GPUs. Our choice of a block-based volume rendering strategy, tailored for resource-independent web environments, achieves seamless integration between blocks. Our carefully designed LOD generation and refinement strategy support dynamic loading, minimizing necessary resources on the web while ensuring the best visual experience. Extensive experiments have also fully proved the effectiveness of City-on-Web.
A Derivation of Block-based Volume Rendering
Traditional volume rendering methods, like those used in Block-NeRF [35] and Mega-NeRF [39], necessitate accessing radiance and opacity for all sampling points, even if the points belong to different blocks. This requirement means that when the rendering resources of multiple shaders cannot communicate with each other, their volume rendering algorithm cannot work. On the contrary, our block-based volume rendering algorithm is tailored for such resource-independent settings, allowing for the independent rendering of each block before blending their outputs. In this section, we will prove the equivalence between this resource-independent volume rendering and the traditional volume rendering under Lambertian conditions.
In traditional volume rendering, consider a ray $\boldsymbol{r}$ with $M$ sampling points. For each sampling point $p_{i}$, the diffuse color, feature, and opacity are represented as $\boldsymbol{c}_{i}$, $\boldsymbol{f}_{i}$, and $\alpha_{i}$ respectively. The final diffuse color $\boldsymbol{C}_{d}$ of the ray $\boldsymbol{r}$ can be obtained by
$\hat{\boldsymbol{C}}_{d}(\boldsymbol{r})=\sum_{i=1}^{M}\prod^{i-1}_{j=1}(1-\alpha_{j})\alpha_{i}\boldsymbol{c}_{i}$ (9)
In block-based volume rendering, let’s assume the ray, with $M$ sampling points, traverses through K blocks. Within block $\mathcal{B}_{k}$, there are $N_{k}$ sampling points. For sampling point $p_{i}^{k}$ in block $\mathcal{B}_{k}$, the diffuse color, feature, and opacity output by the shader or submodel of block $\mathcal{B}_{k}$ are represented as $\boldsymbol{c}^{k}_{i}$, $\boldsymbol{f}^{k}_{i}$, and $\alpha^{k}_{i}$, respectively.
$\boldsymbol{c}^{k}_{i}=\boldsymbol{c}_{N_{1}+\cdots+N_{k-1}+i}$ (10)
$\boldsymbol{f}^{k}_{i}=\boldsymbol{f}_{N_{1}+\cdots+N_{k-1}+i}$
$\alpha^{k}_{i}=\alpha_{N_{1}+\cdots+N_{k-1}+i}$
Then, for this ray, diffuse color $\boldsymbol{C}^{k}_{d}$, specular feature $\boldsymbol{F}^{k}$ and opacity $\alpha^{k}$ of block $\mathcal{B}_{k}$ are caculated as follows:
$\boldsymbol{C}^{k}_{d}$ $=\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{c}_{i}^{k}$ (11)
$\boldsymbol{F}^{k}$ $=\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{f}_{i}^{k}$